OpenStudy (anonymous):

what is the equation of the quadratic graph with a focus of (8,-8) and a directrix of y=-6

OpenStudy (anonymous):

A. F( x)= -1/4(x-7)+1 B.F(x)=-1/4(x-8)^2 C.F(x)=-1/4(x-8)^2-7 D.F(x)= 1/4(x-7)^2